

********************************************************************
UARS WEEKLY STATUS REPORT - 12 AUG 2005
********************************************************************
***********
SOLSTICE
***********
Nothing received.
***********
SUSIM
***********
SUSIM's final data was gathered on August 1. No more operations are
planned. We are preparing our (perhaps) final data version.
***********
PEM
***********
HEPSA data files are version 2. Currently, HEPSA data between 1991274
and 2004273 has been reprocessed and archived at the GSFC DAAC. HEPSA data
between 2004274 and 2005099 has been processed and is waiting for archival
at the DAAC.
MEPS data files are version 1. Processed MEPS data files at the DAAC
archive are between 1991274 and 2004259. MEPS data between 2004260 and
2005099 has been processed and is waiting for DAAC archival.
HEPSB data files are version 1. HEPSB processing code has past testing
and has been submitted to the project for approval.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered since updates have occurred.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html. The additional electron data for
the October and November storms of 2003 has been moved and is now located at
http://climatology.space.swri.edu/uars_fromRudy.html. For those users of
SDDAS, update your UARS Meta data and these data times will appear
with full functionality. Proton data from these storms are now located at
http://climatology.space.swri.edu/UARS_images_from_Rudy_(Apr._2004).html
Requests have been received for storm spectral data. Overview plots were
produced and these plots may be found on the web at
http://climatology.space.swri.edu/UARS_images_postedJune2004.html
for the storms of October 29 of 1991, November 8-9 of 1991, April 6-7 of 2000,
March 31 - April 1 of 2001, April 11-12 of 2001, November 6-7 of 2001, and
November 11-12 of 2001.
PEM has received a request for data from the super storm of November
2004. The Dst value showed sudden commencement early on November 7, 2004
reaching a Dst value of about -380 gamma early on November 8, 2004.
PEM was not active at that time and became active late on November 8, 2004
in day only mode. A second peak of about -280 gamma was observed on November
11, 2004. PEM became active in day-light only mode late on November 8. PEM
survey spectrograms for November 8-18 of 2004 may be found at:
http://climatology.space.swri.edu/UARS_images_fromRudy.html
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations until the end-of-mission.
Daily processing of instrument Daily Activity Plans continues
with no problems for HALOE and PEM (PEM is to be used for power
management purposes, on as needed basis).
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 30 July 2005 (Orbit 75993, GMT Day 211) through 05
August 2005 (Orbit 76099, GMT Day 217).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
SUSIM 213/2132 76037 POWER OFF
SOLSTICE 213/2213 76037 DAY-ONLY TO OFF
PEM 215/1223 76061 OFF TO DAY-ONLY
PEM 215/1448 76062 DAY-ONLY TO OFF
SOLSTICE 215/1808 76065 POWER ON
SOLSTICE 215/2025 76066 POWER OFF
The beta angle ranged between the angles of 33.9 degrees to 15.6
degrees for this report period. The beta angle is now decreasing
towards a minimum angle of 4 degrees on 09 August 2005 (DOY 221).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 7.60 to 7.20 degrees C. Additional
battery performance data for each battery on 29 July 2005 and 05
August 2005 are provided in Appendix D.
The clock error ranged from 10.2 to -10.2 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 07
SUSIM = 01
SOLSTICE = 03
PEM = 02
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 02
Two (2) Anomaly Reports remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
Two (2) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The UARS OBC clock rollover was successfully completed as scheduled on
02 August 2005 during the 1409Z contact. The epoch date was changed
from 31 December 1996 to 31 December 2004. The FOT would like to
thank all entities involved in doing a great job preparing for and
executing the clock jump.
This week the FOT continued to closely monitor the batteries. The
differential voltage on battery 2 has been consistently in the -400mV
range this week. Also the max temperature on battery 2 has remained
at a high of 11.76 C during this time frame. Nominally during times
of high differential voltages, a higher VT level is commanded, but the
high temperature on battery 2 suggests that we drop to a lower VT
level. NASA power branch recommends that we stay at VT6 and only make
a change if the differential on battery 2 goes higher. UARS will look
at using VT7 for the yaw if conditions dictate a change is needed.
Dimitrios Mantziaras has accepted another position within Honeywell.
Effective 08 August 2005, Mehul Patel will be the UARS Acting Lead
Engineer. Mehul will coordinate all day-to-day operations, phone
calls, instrument operations, yaw maneuver planning, etc. Dimitrios
will be available for UARS consultation, and will be the Lead Engineer
for De-Boost activities.
MEETINGS
The FOT attended a meeting with the NASA power branch to discuss the
high differential voltages being seen on battery 2 during the end of
night. Also discussed were the power plans for the upcoming yaw on 09
August 2005.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 09 August 2005 (DOY 221).
This maneuver will be a Reverse to Forward Flight maneuver (Southern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 05 August 2005
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
76004 211/1642 05-237 MPS VT L6 TO L5 // BATSOC
76008 211/2320 05-238 MPS VT L5 TO L6
76037 213/2132 05-239 SUSIM POWER OFF
76037 213/2213 05-240 SOLSTICE DAY-ONLY TO OFF
76047 214/1307 05-241 OBC PRE OBC EPOCH CLOCK RESET
76048 214/1451 05-242 OBC OBC EPOCH CLOCK RESET
76051 214/1948 05-243 OBC FSW VER CHANGE TO A & BANK DUMPS
76051 214/2028 05-243 OBC FSW VER CHANGE TO A & BANK DUMPS
76061 215/1223 05-244 PEM PEM OFF TO DAY-ONLY
76062 215/1448 05-245 PEM PEM DAY-ONLY TO OFF
76065 215/1808 05-246 SOLSTICE POWER ON
76066 215/2025 05-247 SOLSTICE POWER OFF
76077 216/1325 05-248 OBC/CLK CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
76096 217/1938 05-249 OBC/CLK CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 05 August 2005
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 05 August 2005
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
76077 216/1325 CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
76096 217/1943 CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 05 August 2005
29 July 2005 - GMT Day 209 Beta = 34.0 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 336.0/ 224.0 -2.42/-3.68 +00.0/+00.0 05.4 0.00 0.0
2 -117.6/-268.8 +10.73/+9.72 +28.8/-16.0 26.7 1.09 19.9
3 22.4/-677.6 +4.50/+3.87 +19.2/-10.4 26.7 1.06 13.3
05 August 2005 - GMT Day 217 Beta = 15.6 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 728.0/ 728.0 -2.42/- 3.68 +00.0/+00.0 04.8 0.00 0.0
2 -112.0/-425.6 +12.10/+11.07 +32.8/-17.6 25.92 1.08 21.7
3 44.8/-700.0 +4.50/+ 3.87 +19.2/-10.4 25.76 1.05 12.7
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
76004 211/1642 LV6 -> LV5
76008 211/2320 LV5 -> LV6
APPENDIX E
UARS ANOMALY REPORTS
Week ending 05 August 2005
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-032 FOT 76044 214/0831 30832 Generic Late Acquisition
05-038 FOT 76062 215/1405 40360 Dropped frames during R/T
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-036 FOT 75831 200/0440 N/A ECLIPSE Command Hung-up
05-037 FOT 75859 202/0216 40172 Late Acq: bad weather @ Guam
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-032 FOT See List July 05 30832 Generic Late Acquisition
05-038 FOT 76062 215/1405 40360 Dropped frames during R/T